Evaluate: 3^-2 ??????????????????????????????????????????????

 Evaluate:
3^-2 ??????????????????????????????????????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

3-^2= (1/3)^2
[ a^-m = 1/a^m ]
= 1/9

Thanks